find the value for m for which equations have equal roots                       x2-15-m(2x-8)=0
 

Grade:12

1 Answers

Nishant Vora IIT Patna
8 years ago
Dear student,
Its quadratic equation
if two roots are equal that mean D=0
==> b2 – 4ac =0
x^2-15-m(2x-8)=0
x^2-2mx + (8m - 15)=0
So
(-2m)^2 -4(8m - 15) =0
Solve this now